# Fuzzy to Clear: Elucidating the Threat Hunter Cognitive Process and Cognitive Support Needs

### v2

@@ 3.2 Data Analysis
Throughout our data analysis process, we documented the steps and decisions made, ensuring transparency in the activities performed. A high-level overview of this process is presented in the following subsections. Focusing on key insights, we present our three rounds of data analysis. The first round entails a thematic analysis [49] of the observation sessions (see Fig. 1-c). Building on top of the observations notes and results that emerged during Round 1, two more rounds of data analysis were conducted: Round 2, quantitative analysis (see Fig. 1-g), and Round 3, identification of areas with problems and opportunities for improvement (see Fig. 1-i).
@@ 3.2.1 Round 1: Thematic data analysis
First, each researcher translated their notes from the observation sessions into “cards” within an online collaborative board, Miro [^5] Miro: https://miro.com/. The cards were then associated with “categories” to roughly sort the cards into groups. These activities were done by each researcher individually (corresponding to Steps 1 and 2 in Fig. 2).
@@
It is worth noting that while we employed an inductive approach to analyze the data collected (from the observation notes to the themes), the initial set of categories associated to the cards was derived from the core information concepts identified in our previous study within the threat hunting context ([40]). In the Appendix, we present further details of our coding scheme, including the list of categories (Table V), examples of associated observation notes (cards), and the list of themes (Table VII).
@@ 10 Sharing_Report 14
Figure 3: The threat hunter process of building a mental model during a threat hunting session. A total of six intentions (orange boxes on the bottom) and six activities (blue boxes) are organized in three phases (grey boxes).
@@ 5 Building a Mental Model in a Threat Hunting Session
II b. Explore Searching / Exploring
c. Enrich Annotating
@@
III e. Elucidate Documentation / Reports
@@
To provide a walk-through of the threat hunter process of building a mental model, we illustrate a threat hunting session performed by a TH named Olivia [^6] Olivia is inspired by the persona reported in [40].. Olivia is an experienced TH working as an external consultant who uses the same SIEM tool as her clients. Olivia is also the team lead of a group of THs who share the same portfolio of clients.
@@
We consider our threat hunter process of building a mental model representation to be abstract but detailed enough to help inform strategies and inspire innovation to support the THs. We next present the suggested five design propositions that emerged from our study in Section 6.
@@ 6 Design Propositions
This section introduces five design propositions (DP) that emerged from our comprehensive data analysis (explained in Section 3). These propositions, informed by our experience as tool designers, aim to address the intentions of the threat hunter process of building a mental model (see Figure 1-k). The five propositions we propose include:
@@
In the next subsections, the five DP are presented at a high level but with what we believe to be sufficient detail to convey their potential impact. For each DP, we provide the goal, problems, potential solutions (or features), and user stories. For the examples of user stories, we refer to the TH personas created by Hill et al. [40]: Olivia (the creative team lead TH who hunts proactively and prides herself in her leadership skills), Jay (the newer TH, fresh from an academic track with excellent analytic skills and a reactive hunting style), and Thomas (the most experienced TH that works in a small team with an intuitive hunting style).
It is important to note that aspects regarding usability or friction points in the current UI of the TH tool were not the target of our study. Still, some notes can be observed about the UI friction point in the presentation of the findings in Section 4. Our study is also not intended to be an exhaustive or definitive set of DP based on our observations, analysis, and discussions. Therefore, we selected the five prominent DP to inspire continued exploration. We are confident these DP provide a foundation for innovation and advancement to support the threat hunting cognitive process.
@@ 6.1 Creating a Story or Timeline of Events
DP 1 Creating a Story or Timeline of Events aims to support the TH’s process of formulating a clear story of what’s happened from an initially fuzzy state, track progress, initialize, store and update their mental model, and facilitate collaboration.
Related problems include: (a) many disconnected threads of information to keep track of; (b) high cognitive load while hunting; and (c) THs rely on their memory to store information.
Potential solutions should consider:
  • A mind map integrated into the TH tool.
  • A timer to know how long a hunt has taken.
  • Percentage coverage indicator to track progress.
@@
  • The ability to collaborate with peers on a hunt.
  • Collect information and store thoughts in story frames buckets (abstract container) with notes and annotations.
  • Find information quickly with search filters.
User stories (examples for DP 1):
  • As Olivia, I want to externalize (draw) the story of the hunt I’m working on so that I can clarify my thoughts and reduce my cognitive load.
@@
  • As Jay, I want to have a way to save interesting items and sort items into abstract containers so I can review the items easily and begin to clarify my thoughts.
  • As Thomas, I want to add events, machines, and users directly from my hunting tool to my externalized mental model so that my externalization has backlinks to the data and includes relevant context.
  • As Thomas, I want to capture the temporal dimension of my externalized mental model so that I can describe the influence and spread of suspicious activity.
  • As Olivia, I want to use my externalized mental model as a starting point to generate (elucidate) formal reports for clients and management.
  • As Jay, I want my externalized mental model to be used by my hunting tool to assist in exploration (e.g., suggestions, in context analytics, or shortcut for searching).
  • As Olivia, I want my threat hunting tool to suggest whether or not to stop hunting based on the findings and comprehensiveness of my externalized mental model so that I can conclude hunts with confidence and avoid ending too early (potentially missing evidence), or ending too late (and taking time away from other hunts).
  • As Jay, I want to be able to save and use a previous workflow for a new hunt.
  • As Thomas, when I start my hunt I want to be able to examine the stories from the previous shift to be able to orient myself in the current threat situation.
@@ 6.2 Visualizing and Navigating Connections
DP 2 Visualizing / Navigating Connections (Spatial) goal is to orient the TH to the map of the client’s system/network, help THs navigate to related connections while maintaining bearings (orientation) and let THs be able to sense of where they are within the map.
@@
Potential solutions should consider:
@@
User stories (examples for DP 2):
@@
  • As Olivia, I want to explore the event details of several different events without losing track of where I am, so that I can synthesize and make sense of what’s happening.
  • As Thomas, I want to see what happened before and after an event, so that I can identify correlations of events related in time.
@@ 6.3 Creatively Expanded Search
DP 3 Creatively Expanded Search goal refers to patterns in how THs seek information and the possibility of constructing an extended search capability that matches the direction of inquiry.
Related problems include: (a) TH wants to see if there are similar execution patterns happening across a variety of machines; and (b) TH wants a way to quickly filter subsets of data (based on a variety of different things).
Potential solutions should consider:
  • “If I had regex…” is a thing to imagine, but maybe not the thing to build, so what are some creative alternatives that can still support pathways of inquiry and exploration? Some options to explore could be: clustering / fingerprint similar patterns; regex; filtering; sorting.
@@ 6.4 Waypoints and Note-Taking
DP 4 Waypoints and Note-Taking refers to TH be able to add metadata and notes that support the construction of the story of the threat that is unfolding in their mental model.
Related problems include: (a) TH needs the ability to take more notes within the TH tool; (b) TH loses notes or bookmarks during the threat hunting; (c) TH misses what other THs did during the same investigation; and (d) TH difficulty to share notes and bookmarks to inform future hunts.
Potential solutions should consider:
  • Bookmarking of events and entities within the tool to inform future hunts (not only for documentation).
  • Waypoints, include metadata (as one key difference from bookmarks) and are constructed to support navigation as part of a “journey.”
  • Rank or like (or dislike) previous comments to support their agreement with the hypothesis or to say something is no longer critical.
  • Keep notes within main TH tool to future hunts (not just documentation).
  • Create a thread of comments that link through multiple events.
  • Be able to collaborate on creating and using annotations.
  • Create a set of waypoints that trace a path through the data and that can be shared with others.
User stories (examples for DP 4):
@@
  • As Thomas, I want to annotate the event data in my hunting tool with waypoints, so that I can record my path and so that others and I can retrace my steps and ensure the full coverage of the data.
  • As Olivia, I want an integrated view and history of annotations so that I can collaborate more easily with my team and discuss team agreement of the notes.
  • As Olivia, I want to enhance the annotations with a agreement/disagreement feature so that I can collaborate more easily with my team and coordinate different beliefs across the team.
@@ 6.5 Integrate External Resources
Related problems include: (a) context switching caused by internet searches for information supporting their hunting activities; (b) online information resource (e.g., documentation sites) can be difficult to locate or search through; and (c) the first or most obvious search results may be second-hand information (e.g., from users on Reddit, Stack Overflow), which could be poisoned by bad actors.
Potential solutions should consider:
  • Automate the lookup of executable hashes integrated into the tools.
  • Integration of common bookmarked resources.
@@
  • As Olivia, I want to have common, external, and previously bookmarked resources integrated into my threat hunting tool so that I can access these resources in a more streamlined way within my threat hunting process and ensure the information I’m referencing is accurate.
@@ 7.1.1 Empirical evidences applicable to other threat hunting contexts
The profile of the THs we studied includes individuals working in collaborative team environments and consulting for clients within the private sector using the same tools. In addition, THs in our study used the industry partner tool (information omitted for the review process) as their main tool. However, despite our focused context, the findings described in our results apply to other tools used by THs in general (e.g., including Splunk and LogRhythm, which offer SIEM and UEBA capabilities in a single product).
Our findings also resonate closely with established threat hunting models, such as TaHiTI [2], affirming their relevance and potential applicability to the broader threat hunting scope. Unlike TaHiTI, which focuses solely on structured hunting, our proposal accommodates both structured and unstructured hunting workflows. This dual faceted approach aligns with the comprehensive perspective found in earlier references (e.g., [50, 9]), however, our model maintains a balanced level of detail to enhances its usability, relevance, clarity, and applicability.
@@ 7.1.2 Threat hunter cognitive efforts
During our observations, we noticed the significant cognitive effort that THs invest in their memory, underscoring the importance of supporting the construction, externalization, and sharing of a robust mental model. We explicitly incorporated the transition of a TH’s mental model from fuzziness to clarity into our model. This transition highlights the cognitive processes involved, often hidden or underemphasized in other references, even those concentrating on cognitive aspects like Andrade and Yoo [10]. By expanding on the Mental Model, we aim to represent the intricate and indirectly observable processes of information processing that are crucial during threat hunting sessions.
Our model seeks to elucidate these human thought processes without overloading the visual representation with extraneous activities, thus maintaining clarity and focus on core cognitive tasks. Unlike many other models within the security scope that lean towards developing automation and AI, our model emphasizes supporting the human aspects of threat hunting. We aim to capture the intentions and activities performed by THs rather than automating cognitive tasks such as pattern recognition (hence preserving the essence of human cognitive engagement in cybersecurity).
@@ 7.1.3 Bringing attention to new activities
Our proposed threat hunter process of building a mental model strives to balance practicality and applicability to support the design propositions derived from our findings and our prior experience designing tools. It aligns with the practical, applicable nature of TaHiTI model[2], avoiding the overly detailed approach of other references (e.g., [50] or [10]) while also steering clear of the high-level abstraction (e.g., [21]). This balanced approach ensures that our proposed model remains accessible and valuable for practitioners, enabling them to apply our insights effectively in real-world threat hunting scenarios. Considering this and the specific context of our observations (i.e., THs using SIEM and UEBA tools, working with clients using the same tools), we have emphasized two critical activities in our representation: Annotating and Materialize Mental Model. These strategic additions underscore the practical necessity for THs to document their findings and thoughts throughout their investigative processes (Threat Hunting Session).
@@ 7.2.2 Observations of replayed threat hunting sessions
During our observation sessions, the TH participants presented how they worked, shared the tools they used and processes followed, and demonstrated their hunting sessions (including actual client data discussing real scenarios, even though not “live”). The participants selected the content they wanted to share with us as a replay of a previous threat hunting session, with thought and consideration about what they wanted to showcase. Consequently, they might have omitted specific steps or presented their hunting activities in a different light than they would have in real time. This approach provided valuable insights, but we acknowledge that this method might not fully capture the spontaneous and dynamic nature of live threat hunting.
@@ Acknowledgments
We are adding information to support the data analysis process description (coding scheme). Table V shows the list of categories descriptions and examples of the cards associated with them. Table VI shows the three cards with more categories associated with them. Figure 5 shows the categories’ similarity based on the cards related to them. Table VII shows the list of themes (alphabetically ordered) and their brief description—full theme description including an example of observation note was described in Section 4.
@@
TABLE VI: Cards with more categories associated with (total of seven each) and the final theme related to them.

### v3

@@ 3.2 Data Analysis
Throughout our data analysis process, we documented the steps and decisions made, ensuring transparency in the activities performed. A high-level overview of this process is presented in the following subsections. Focusing on key insights, we present our three rounds of data analysis. The first round entails a thematic analysis [50] of the observation sessions (see Fig. 1-c). Building on top of the observations notes and results that emerged during Round 1, two more rounds of data analysis were conducted: Round 2, quantitative perspective (see Fig. 1-g), and Round 3, identification of areas with problems and opportunities for improvement (see Fig. 1-i).
@@ 3.2.1 Round 1: Thematic data analysis
First, each researcher translated their notes from the observation sessions into “cards” within an online collaborative board, Miro. [^5] Miro: https://miro.com/ The cards were then associated with “categories” to roughly sort the cards into groups. These activities were done by each researcher individually (corresponding to Steps 1 and 2 in Fig. 2).
@@
It is worth noting that while we employed an inductive approach to analyze the data collected (from the observation notes to the themes), the initial set of categories associated to the cards was derived from the core information concepts identified in our previous study within the threat hunting context ([18, 51]). In the Appendix, we present further details of our coding scheme, including the list of categories (Table V), examples of associated observation notes (cards), and the list of themes (Table VII).
@@ 10 Sharing_Report 14
Figure 3: The threat hunter process of building a mental model during a threat hunting session. This visual representation includes a total of six intentions—(a) Examine, (b) Explore, (c) Enrich, (d) Externalize, (e) Elucidate, and (f) Engage—and six related activities (blue boxes positioned above the intentions) organized in three phases (grey boxes labelled with Session). We highlight the Mental Model to capture the complex, often unseen, cognitive processes central to threat hunting.
@@ 5 Building a Mental Model in a Threat Hunting Session
II b. Explore Search / Explore
c. Enrich Annotate
@@
III e. Elucidate Document / Report
@@
To provide a walk-through of the threat hunter process of building a mental model, we illustrate a threat hunting session performed by a TH named Olivia. [^6] Olivia is inspired by the persona reported in [18]. Olivia is an experienced TH working as an external consultant who uses the same SIEM tool as her clients. Olivia is also the team lead of a group of THs who share the same portfolio of clients.
@@
We consider our threat hunter process of building a mental model representation to be abstract but detailed enough to help inform strategies and inspire innovation to support the THs. We further discuss the implications of our model in Subsection 7.1. In the next section, we present five design propositions that emerged from our study.
@@ 6 Design Propositions
This section introduces five design propositions (DP) that emerged from our comprehensive data analysis (explained in Section 3). These propositions, informed by our experience as tool designers, aim to support the intentions of the threat hunter’s process of building a mental model (see Figure 1-k). The five propositions we propose include:
@@
In the next subsections, the five DP are presented at a high level but with what we believe to be sufficient detail to convey their potential impact. Although some DPs are more specific or detailed than others, for each DP, we provide the goal, problems, potential solutions (or features), and user stories. For the examples of user stories, we refer to the TH personas created by Hill et al. [18]: Olivia (the creative team lead TH who hunts proactively and prides herself in her leadership skills), Jay (the newer TH, fresh from an academic track with excellent analytic skills and a reactive hunting style), and Thomas (the most experienced TH that works in a small team with an intuitive hunting style).
It is important to note that aspects regarding usability or friction points in the current UI of the TH tool were not the target of our study. Still, some observations were made about the UI friction points in the presentation of the findings in Section 4. Our study is also not intended to be an exhaustive or definitive set of DP based on our observations, analysis, and discussions. Therefore, we selected the five most prominent DP to inspire continued exploration. We are confident these DP provide a foundation for innovation and advancement to support the threat hunting cognitive process.
@@ 6.1 Creating a Story or Timeline of Events
DP 1 Creating a Story or Timeline of Events aims to support the TH’s process of formulating a clear story of what happened during a possible attack, as the TH’s mental model goes from an initially fuzzy state, to a clear understanding, allowing the TH to track progress, update their mental model, and collaborate with their team.
Related problems include: (a) many disconnected threads of information to keep track of; (b) high cognitive load while hunting; and (c) THs relying on their memory to store information.
Potential solutions could consider:
  • A mind map integrated into the TH tool that allows the TH to keep track of emerging connections and findings.
  • A timer to know how long a hunt has taken so far.
  • Percentage coverage indicator to track the progress of what has been reviewed.
@@
  • The ability to share a partial understanding with a mind map to collaborate with peers on a hunt.
  • Collect information leads into story frame buckets (abstract container) with notes and annotations that can be organized into a story.
  • Find information quickly with search filters that are driven by the contents of the story frames.
User stories (a more comprehensive list of user story examples for DP 1 is provided in Appendix-Table VIII):
  • As Olivia, I want to externalize (draw) the story of the hunt I am working on so that I can clarify my thoughts and reduce my cognitive load.
@@ 6.2 Visualizing and Navigating Connections
DP 2 Visualizing / Navigating Connections (Spatial) goal is to orient the TH to the map of the client’s system/network, help THs navigate to related connections while maintaining bearings (orientation), and let THs be able to sense of where they are within the map.
@@
Potential solutions could consider:
@@
User stories (a more comprehensive list of user story examples for DP 2 is provided in Appendix-Table VIII):
@@ 6.3 Creatively Expanded Search
DP 3 Creatively Expanded Search goal refers to patterns in how THs seek information and the possibility of constructing an extended search capability that matches the TH’s intuitive direction of inquiry.
Related problems include: (a) TH wants to see if there are similar execution patterns happening across a variety of machines when seeing a suspicious execution pattern, and have no direct way to support this inquiry; and (b) TH wants a way to quickly filter subsets of data (based on a variety of different things).
Potential solutions could consider:
  • Search pathways that can find events by their similarity to an execution pattern, possibly by fingerprinting or clustering similar patterns up front.
  • The ability to save execution patterns that represent common attack patterns as they are learned.
  • The ability to search across machines by execution/attack pattern.
@@ 6.4 Waypoints and Note-Taking
DP 4 Waypoints and Note-Taking refers to TH being able to add metadata and notes that support the construction of the story of the threat that is unfolding in their mental model.
Related problems include: (a) TH needs the ability to take more notes within the TH tool; (b) TH loses notes or bookmarks during the threat hunting; (c) TH misses what other THs did during the same investigation; and (d) difficulty of sharing notes and bookmarks to inform future hunts.
Potential solutions could consider:
  • Bookmarking events and entities within the tool to inform future hunts (not only for documentation).
  • Supporting waypoints that are constructed during navigation that help the TH visualize their “journey” through the data, and return to previous waypoints.
  • Rank or like (or dislike) previous comments to support their agreement with a hypothesis, or say something is no longer critical.
  • Keeping notes within main TH tool to support future hunts (not just documentation).
  • Supporting a thread of comments that link through multiple events.
  • Being able to collaborate on creating and using annotations.
  • Creating a set of waypoints that trace a path through the data and that can be shared with others.
User stories (a more comprehensive list of user story examples for DP 4 is provided in Appendix-Table VIII):
@@ 6.5 Integrate External Resources
Related problems include: (a) context switching caused by internet searches for information supporting hunting activities; (b) online information resources (e.g., documentation sites) can be difficult to locate or search through; and (c) the first or most obvious search results may be second-hand information (e.g., from users on Reddit, Stack Overflow), which could be poisoned by bad actors.
Potential solutions could consider:
  • Automate the lookup of executable hashes and integrate into the TH tool.
  • Integration of common external bookmarked resources into the TH tool.
@@
  • As Olivia, I want to have common, external, and previously bookmarked resources integrated into my threat hunting tool so that I can access these resources in a more streamlined way within my threat hunting process and ensure the information I am referencing is accurate.
@@ 7.1.1 Empirical evidences applicable to other threat hunting contexts
The profile of the THs we studied includes individuals working in collaborative team environments and consulting for clients within the private sector using the same tools. In addition, THs in our study used the industry partner tool (OpenText, ArcSight Intelligence) as their main tool. However, despite our focused context, the findings described in our results apply to other tools used by THs in general (e.g., including Splunk and LogRhythm, which offer SIEM and UEBA capabilities in a single product).
Our findings also resonate closely with established threat hunting models, such as TaHiTI [2], affirming their relevance and potential applicability to the broader threat hunting scope. Unlike TaHiTI, which focuses solely on structured hunting, our proposal accommodates both structured and unstructured hunting workflows. This dual faceted approach aligns with other empirical descriptive TH processes (e.g., [10, 3]), however, our model maintains a balanced level of detail to enhances its usability, relevance, clarity, and applicability.
@@ 7.1.2 Threat hunter cognitive efforts
During our observations, we noticed the significant cognitive effort that THs invest in their memory, underscoring the importance of supporting the construction, externalization, and sharing of a robust mental model. Considering the clarity of the TH’s mental model directly impacts the effectiveness of reporting, communication, and collaboration with clients and peers, we explicitly incorporated the transition of a TH’s mental model from fuzziness to clarity into our model. This transition highlights the cognitive processes involved, often hidden or underemphasized in other references, even those concentrating on cognitive aspects like Andrade and Yoo [11]. By expanding on the Mental Model, we aim to represent the intricate and indirectly observable processes of information processing that are crucial during threat hunting sessions.
Our model seeks to elucidate these human thought processes without overloading the visual representation with extraneous activities, thus maintaining clarity and focus on core cognitive tasks. Unlike many other models within the security scope that lean towards developing automation and AI, our model emphasizes supporting the human aspects of threat hunting. We aim to capture the intentions and activities performed by THs rather than automating cognitive tasks such as pattern recognition—hence preserving the essence of human cognitive engagement in cybersecurity.
@@ 7.1.3 Bringing attention to new activities
Our proposed threat hunter process of building a mental model strives to balance practicality and applicability to support the design propositions derived from our findings and our prior experience designing tools. It aligns with the practical, applicable nature of TaHiTI model[2], avoiding the overly detailed approach of other references (e.g., [10, 11, 52]) while also steering clear of the high-level abstraction (e.g., [3, 26]). This balanced approach ensures that our proposed model remains accessible and valuable for practitioners, enabling them to apply our insights effectively in real-world threat hunting scenarios. Considering this and the specific context of our observations (i.e., THs using SIEM and UEBA tools, working with clients using the same tools), we have emphasized two critical activities in our representation: Annotate and Materialize Mental Model. These strategic additions underscore the practical necessity for THs to document their findings and thoughts throughout their investigative processes (Threat Hunting Session).
@@ 7.2.2 Observations of replayed threat hunting sessions
During our observation sessions, the TH participants presented how they worked, shared the tools they used and processes followed, and demonstrated their hunting sessions (including actual client data discussing real scenarios, even though not “live”). The participants selected the content they wanted to share with us as a replay of a previous threat hunting session, with thought and consideration about what they wanted to showcase. Consequently, they might have omitted specific steps or presented their hunting activities in a different light than they would have in real time. This observational approach provided valuable insights, revealing unspoken or unconscious practices that would have been difficult to uncover through interviews, as participants may not recall or articulate the full nuance of their activities. However, we acknowledge that replayed sessions may present a more polished or linear narrative than live investigations, introducing a potential narrative bias. To account for this, we triangulated the observations across multiple participants and sessions. Additionally, our model explicitly emphasizes the iterative and non-linear nature of threat hunting, reflecting how activities such as Search / Explore, Annotate, and Materialize Mental Model can recur and overlap rather than unfold sequentially.
@@ Acknowledgments
We are adding information to support the data analysis process description (coding scheme). Table V shows the list of categories descriptions and examples of the cards associated with them. Table VI shows the three cards with more categories associated with them. Figure 5 shows the categories’ similarity based on the cards related to them. Table VII shows the list of themes (alphabetically ordered) and their brief description—full theme description including an example of observation note was described in Section 4. Finally, Table VIII shows user story examples for each Design Proposition—presented in Section 6.
@@
TABLE VI: Cards with more categories associated with (total of seven each) and the final theme related to them
@@
TABLE VIII: Examples of user stories for each design proposition
Design Proposition User Story
DP1. Creating a Story or Timeline of Events As Olivia, I want to externalize (draw) the story of the hunt I am working on so that I can clarify my thoughts and reduce my cognitive load. As Thomas, I want to add notes to my externalized mental model so that I can explain the data to myself and others. As Olivia, I want to share my externalized mental model with other THs to align our mental models and share our findings during an active hunt. As Jay, I want to have a way to save interesting items and sort items into abstract containers so I can review the items easily and begin to clarify my thoughts. As Thomas, I want to add events, machines, and users directly from my hunting tool to my externalized mental model so that my externalization has backlinks to the data and includes relevant context. As Thomas, I want to capture the temporal dimension of my externalized mental model so that I can describe the influence and spread of suspicious activity. As Olivia, I want to use my externalized mental model as a starting point to generate (elucidate) formal reports for clients and management. As Jay, I want my externalized mental model to be used by my hunting tool to assist in exploration (e.g., suggestions, in context analytics, or shortcut for searching). As Olivia, I want my threat hunting tool to suggest whether or not to stop hunting based on the findings and comprehensiveness of my externalized mental model so that I can conclude hunts with confidence and avoid ending too early (potentially missing evidence), or ending too late (and taking time away from other hunts). As Jay, I want to be able to save and use a previous workflow for a new hunt. As Thomas, when I start my hunt I want to be able to examine the stories from the previous shift to be able to orient myself in the current threat situation.
DP2. Visualizing / Navigating Connections (Spatial) As Thomas, I want to visualize and see where I am within a spatial map of the client’s system/network so that I can orient more easily to understand the activity I’m seeing. As Jay, I want to navigate from one entity to other related entities based on interactions in the data, so that I can explore the connections and make sense of what’s happening. As Olivia, I want to explore the event details of several different events without losing track of where I am, so that I can synthesize and make sense of what’s happening. As Thomas, I want to see what happened before and after an event, so that I can identify correlations of events related in time.
DP3. Creatively Expanded Search As Jay, I want to search for patterns that are similar to known attack strategies and be able to “find similar” patterns to one I’m seeing. As Thomas, I want to search for patterns in the data that are similar to observed patterns in the attacker’s activities so far, so that I can discover the full scope of compromised machines.
DP4. Waypoints and Note-Taking As Olivia, I want to keep notes within my threat-hunting tool linking together a set of data, so that I can use the notes to help me construct a mental model and to inform future hunts. As Jay, I want to have alternative ways to filter data by annotations (metadata), to facilitate reviewing and working on annotations created by myself or other THs during an investigation. As Thomas, I want to annotate the event data in my hunting tool with waypoints, so that I can record my path and so that others and I can retrace my steps and ensure the full coverage of the data. As Olivia, I want an integrated view and history of annotations so that I can collaborate more easily with my team and discuss team agreement of the notes. As Olivia, I want to enhance the annotations with a agreement/disagreement feature so that I can collaborate more easily with my team and coordinate different beliefs across the team.
DP5. Integrating External Resources As Olivia, I want to have common, external, and previously bookmarked resources integrated into my threat hunting tool so that I can access these resources in a more streamlined way within my threat hunting process and ensure the information I am referencing is accurate. As Jay, I want to have an automated lookup of executable hashes, so I can know immediately if the executable is custom or known.